# Measuring Faithfulness of Chains of Thought by Unlearning Reasoning Steps

### v1

@@ Abstract
When prompted to think step-by-step, language models (LMs) produce a chain of thought (CoT), a sequence of reasoning steps that the model supposedly used to produce its prediction. However, despite much work on CoT prompting, it is unclear if CoT reasoning is faithful to the models’ parameteric beliefs. We introduce a framework for measuring parametric faithfulness of generated reasoning, and propose Faithfulness by Unlearning Reasoning steps (fur), an instance of this framework. fur erases information contained in reasoning steps from model parameters. We perform experiments unlearning CoTs of four LMs prompted on four multi-choice question answering (MCQA) datasets. Our experiments show that fur is frequently able to change the underlying models’ prediction by unlearning key steps, indicating when a CoT is parametrically faithful. Further analysis shows that CoTs generated by models post-unlearning support different answers, hinting at a deeper effect of unlearning. Importantly, CoT steps identified as important by fur do not align well with human notions of plausbility, emphasizing the need for specialized alignment. [^1] Code available at https://github.com/technion-cs-nlp/parametric-faithfulness
Measuring Faithfulness of Chains of Thought
by Unlearning Reasoning Steps
@@ 1 Introduction
Figure 1: An illustration of pff and fur. In order to produce a parameter intervention, we first prompt the model $\mathcal{M}$ to produce an answer $y$ and reasoning chain (CoT). We then segment the reasoning chain and unlearn a single reasoning step from the model. The unlearned model $\mathcal{M}^{*}{}$ is then prompted to produce an answer $y^{*}$. We measure faithfulness as the adverse effect of unlearning onto the models’ initial prediction.
@@
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024). This setup is inherently flawed, as erasing steps from context does not remove knowledge from model parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
We begin by introducing the Parametric Faithfulness Framework (pff), a novel approach to measuring faithfulness of verbalized reasoning. We define necessary components of instances of such a framework in two stages: (1) an intervention on the model parameters; and (2) evaluating parametric faithfulness. See components in Figure 1. pff is a general framework that can be applied to different types of CoT and other free-text explanations.
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), an unlearning-based approach to assessing CoT faithfulness. fur erases information encoded in the CoT from model parameters and assesses whether it affects the model prediction. Concretely, we use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method as the intervention and propose two metrics of quantifying faithfulness of reasoning steps. ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (b) segment a CoT into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure 1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
Through experimental evaluation on four LMs and four MCQA text reasoning datasets, we show we are able to perform valid interventions that successfully affect model predictions while not damaging the models’ general capabilities. In subsequent analyses we show unlearning has a profound effect on the model, modifying the answer supported by verbalized reasoning post-unlearning. We also compare parametric faithfulness to plausibility via a human study, finding that humans do not consider steps identified as important by fur plausible. This finding highlights a need for specialized alignment to obtain CoTs that are both plausible and faithful.
@@ 2 Background and Related Work
Issues with NLE.
Natural language explanations such as CoTs exhibit a number of issues. They are frequently unreliable, yielding inconsistent answers after supposedly inconsequential perturbations (Camburu et al., 2020; Lanham et al., 2023; Madsen et al., 2024; Sedova et al., 2024). Explanations provided by LMs can be non-causal Bao et al. (2025), not aligning with the generated answers. They are often not useful to humans (Joshi et al., 2023) and can contain factually incorrect or hallucinated information (Kim et al., 2021, 2023; Zheng et al., 2023b; Peng et al., 2023; Zhang et al., 2024a). Most importantly, CoTs have been shown to misrepresent the true reasoning process of the LM (Turpin et al., 2023; Roger and Greenblatt, 2023). Turpin et al. show that LMs predictions can be biased by contextual shortcuts, the influence of which is not disclosed in the CoT. In this work, we focus on verifying whether CoTs generated by LMs reflect their parametric beliefs, that is, if the generated reasoning chain is faithful with respect to the model parameters.
Figure 2: The distinction between contextual and parametric faithfulness. Contextual faithfulness measures the effect of context perturbations on the prediction, while parametric faithfulness measures whether verbalized reasoning corresponds to latent reasoning.
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2024), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account for possible confounders, we only use information from generated CoTs to guide unlearning while generating predictions directly by no-CoT prompting, adequately disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation. Experiments show that such structural causal models are rarely implemented by LMs (Bao et al., 2025), confounding the conclusions drawn from contextual faithfulness methods. In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction, assessing parametric faithfulness of individual causal links. The closest to ours is the contemporaneous work of Yeo et al. (2024) which uses activation patching to measure causal effect of corrupting certain hidden states.
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model is still able to retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. See Figure 2 for a visualization of this distinction. Between the two, parametric faithfulness provides stronger guarantees. Models can recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context in place of its obfuscated true reasoning. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters on the output. In other words, to measure parametric faithfulness, we have to intervene on model parameters.
@@
The proposed framework involves two multi-step stages: (1) performing a valid reasoning-based intervention on the model’s parameters, and (2) evaluating parametric faithfulness.
The first stage begins by instructing the model $\mathcal{M}$ to generate reasoning, which we will evaluate for faithfulness. The reasoning steps are used to guide an intervention on $\mathcal{M}$’s parameters, targeting those where a step’s information is stored. This produces a modified model, $\mathcal{M}^{*}$. Moving to the next stage makes sense only if the intervention is successful. Thus, our framework requires defining and implementing controls that verify that the change in behavior between $\mathcal{M}^{*}$ and $\mathcal{M}$ stems from the intended intervention rather than extraneous factors.
In the second stage, faithfulness is assessed with at least one of two evaluation protocols: (1) Instruct both $\mathcal{M}^{*}$ and $\mathcal{M}$ to directly give answers, then compute how often their answers differ. (2) Instruct $\mathcal{M}^{*}$ and $\mathcal{M}$ to reason-then-answer, then compute not only how often their answers differ, but also how often they present different reasoning. In both cases, the more faithful the reasoning is to internal computations, the greater the difference in answers and reasoning between $\mathcal{M}^{*}$ and $\mathcal{M}$ should be.
@@ 4.1 Parameter Intervention
To erase knowledge contained in the verbalized reasoning steps, we use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method. We opt for the KL-divergence regularized variant of NPO. When unlearning, we only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). [^2] We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We conducted experiments with NPO-grad-diff, but results were always slightly worse to NPO-KL. We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We detail hyperparameters in Appendix C.
@@ 4.2 Controls
$S=\frac{1}{\left|\mathcal{D}_{s}\right|}\sum_{k=1}^{\left|\mathcal{D}_{s}\right|}\mathds{1}[y_{k}=y^{*}_{k}],$ (3)
@@ 4.3 Faithfulness Measurements
where $r_{i}$ is the $i$-th reasoning step and $y^{(i)^{*}}$ the prediction made by $\mathcal{M}^{(i)^{*}}$ (after the $i$-th reasoning step is unlearned). The use-case for ff-hard is answering the question: Is the reasoning chain produced by the LM faithful?
@@ 5 Experimental Setup
Base ARC-Challenge OpenbookQA Sports StrategyQA
Model Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen
LLaMA-8B $63.9$ $43.2$ $98.3$ $63.8$ $44.1$ $97.7$ $63.8$ $20.8$ $98.1$ $63.8$ $48.3$ $95.7$ $63.8$
LLaMA-3B $60.4$ $30.7$ $98.1$ $60.2$ $36.6$ $96.1$ $60.2$ $29.3$ $96.6$ $60.3$ $36.3$ $96.9$ $60.3$
Mistral-2 $59.0$ $71.5$ $96.4$ $58.9$ $72.1$ $97.6$ $58.8$ $50.6$ $94.8$ $59.0$ $65.4$ $96.3$ $59.0$
Phi-3 $69.9$ $40.8$ $99.5$ $69.6$ $44.2$ $99.4$ $69.6$ $31.1$ $97.0$ $69.9$ $18.7$ $98.2$ $69.9$
@@
We employ four diverse MCQA datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (ARC-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). We choose MCQA as the target task due to availability of alternative answers, which simplify analysis of how the models’ predictive distribution shifts after unlearning.
To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of up to $250$ instances from the test split of each dataset to balance the question sources. [^4] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix A.
@@
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023); Bentham et al. (2024), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently, the model is prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each model and instance pair. For the prompts used, see Appendix B.
@@
In order to obtain fine-grained information on faithfulness of individual steps, we segment each chain-of-thought into sentences using NLTK (Bird, 2006). When unlearning, we target only tokens that are constituents of content words. [^5] Concretely, we select noun, proper noun, verb, adjective, and number tokens, after running part-of-speech tagging with SpaCy en_core_web_sm (https://spacy.io/). We opt for this approach so as to not unlearn the capability to verbalize reasoning from the models, but only knowledge within the steps, which is a phenomenon we frequently observed prior to making this modification.
Unlearning CoT Steps.
As mentioned in §4.1, we unlearn each CoT step for $5$ iterations. We propagate unlearning loss only from tokens corresponding to content words, target only FF2 layers of the Transformer MLPs, and consider only CoT steps with at least $2$ content tokens. NPO-KL uses a retain set to minimize the KL-divergence between the base and unlearned model’s output distribution and ensure fluency. We sample $4$ CoT steps from other instances as the retain set. Details in Appendix C.
ARC-challenge OpenbookQA Sports StrategyQA
Model fur Add-mistake fur Add-mistake fur Add-mistake fur Add-mistake
LLaMA-8B $\mathbf{39.58}$ $16.15$ $\mathbf{44.33}$ $18.04$ $29.31$ $29.89$ $30.65$ $\mathbf{32.26}$
LLaMA-3B $\mathbf{64.41}$ $31.07$ $\mathbf{68.60}$ $45.93$ $64.88$ $65.48$ $\mathbf{71.02}$ $48.30$
Mistral-2 $\mathbf{40.00}$ $31.58$ $\mathbf{60.00}$ $35.68$ $\mathbf{45.26}$ $36.84$ $\mathbf{48.19}$ $30.21$
Phi-3 $\mathbf{39.05}$ $27.62$ $\mathbf{46.15}$ $38.46$ $\mathbf{53.99}$ $52.15$ $22.22$ $\mathbf{49.74}$

### v2

# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

@@ Abstract
When prompted to think step-by-step, language models (LMs) produce a chain of thought (CoT), a sequence of reasoning steps that the model supposedly used to produce its prediction. Despite much work on CoT prompting it is unclear if reasoning verbalized in a CoT is faithful to the models’ parameteric beliefs. We introduce a framework for measuring parametric faithfulness of generated reasoning, and propose Faithfulness by Unlearning Reasoning steps (fur), an instance of this framework. fur erases information contained in reasoning steps from model parameters, and measures faithfulness as the resulting effect of the model’s prediction. Our experiments with four LMs and five multi-hop multi-choice question answering (MCQA) datasets show that fur is frequently able to precisely change the underlying models’ prediction for a given instance by unlearning key steps, indicating when a CoT is parametrically faithful. Further analysis shows that CoTs generated by models post-unlearning support different answers, hinting at a deeper effect of unlearning. [^1] Code available at https://github.com/technion-cs-nlp/parametric-faithfulness.
Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps
@@ 1 Introduction
Figure 1: An illustration of fur. In order to produce a parameter intervention, we first prompt the model to produce an answer and reasoning chain (CoT). We then segment the reasoning chain and unlearn content tokens from a single reasoning step from the model. The unlearned model is then prompted to produce an answer. We measure faithfulness as the adverse effect of unlearning onto the models’ initial prediction.
@@
Various works set out to explore CoT faithfulness by perturbing tokens within the CoT and observing whether the contextual corruptions affect model prediction (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024). This setup is inherently imprecise, as erasing steps from context does not remove knowledge from model parameters, and the model may still be able to reconstruct corrupted information when generating a prediction. Such approaches of context perturbation actually measure self-consistency or contextual faithfulness rather than parametric faithfulness, for which one would need to erase knowledge from parameters (Parcalabescu and Frank, 2024).
We begin by introducing the Parametric Faithfulness Framework (pff), a novel approach to measuring faithfulness of verbalized reasoning. We define necessary components of instances of such a framework in two stages: (1) an intervention on the model parameters; and (2) evaluating parametric faithfulness. See components in Figure 1. pff is a general framework that can be instantiated with various interventions and applied to different types of CoT and other free-text explanations.
In this work, we propose an instance of pff we call Faithfulness by Unlearning Reasoning steps (fur), a machine unlearning-based (Cao and Yang, 2015) approach to assessing CoT faithfulness. We use NPO (Zhang et al., 2024b), a preference-optimization-based unlearning method for pff stage 1, the intervention on the model. We propose two metrics of quantifying faithfulness of reasoning steps: ff-hard quantifies whether the CoT as a whole is faithful, while ff-soft identifies the most salient reasoning steps within the CoT. Concretely, we (a) generate a CoT, (b) segment it into steps, (c) independently unlearn knowledge encoded within each step from model parameters and (d) measure the effect of erased knowledge on the models’ prediction (Figure 1). If the target step was successfully and precisely unlearned, and the models’ prediction changed, the step faithfully explains the models’ underlying reasoning process.
Through experimental evaluation on four LMs and five MCQA multi-hop reasoning datasets, we show we are able to perform valid interventions affect model predictions while not damaging the models’ general capabilities. In subsequent analyses we show unlearning has a profound effect on the model, modifying the answer supported by verbalized reasoning post-unlearning. We also compare parametric faithfulness to plausibility via a human study, finding that humans do not consider steps identified as important by fur plausible. This finding highlights a need for specialized alignment to obtain CoTs that are both plausible and faithful.
@@ 2 Background and Related Work
Issues with CoTs.
Natural language explanations such as CoTs exhibit a number of issues. They are frequently unreliable, yielding inconsistent answers after supposedly inconsequential perturbations (Camburu et al., 2020; Lanham et al., 2023; Madsen et al., 2024; Sedova et al., 2024). CoTs have been shown to not align with generated answers Bao et al. (2025), they are often not useful to humans (Joshi et al., 2023) and can contain factually incorrect or hallucinated information (Kim et al., 2021, 2023; Zheng et al., 2023b; Peng et al., 2023; Zhang et al., 2024a). Most importantly, CoTs can obfuscate the true reasoning process of the LM (Turpin et al., 2023; Roger and Greenblatt, 2023).
@@
Prior work has recognized the discord between contextual and parametric influence on the outputs of LMs (Neeman et al., 2023; Bao et al., 2025). Prompting models with hypothetical or factually incorrect information causes them to change their otherwise consistently correct predictions (Kim et al., 2021, 2023; Simhi et al., 2024; Minder et al., 2024), highlighting their high sensitivity to context tokens and confounding any conclusions drawn from contextual perturbations applied to reasoning steps. The main issue with work investigating self-consistency is the possibility of the LM reconstructing information obfuscated by the contextual perturbation—despite the verbalized knowledge missing, this reasoning could still be retrieved from the latent space (Yang et al., 2024; Deng et al., 2024). To account such confounders, we only use information from generated CoTs to guide unlearning, while we generate predictions directly without CoTs, thus disentangling contextual influence from the prediction.
@@
Various tests and metrics for quantifying faithfulness of free-text explanations in LMs have previously been proposed (Lanham et al., 2023; Bentham et al., 2024; Atanasova et al., 2023; Siegel et al., 2024). By measuring properties such as sufficiency through simulatability or counterfactual interventions (Atanasova et al., 2023; Lanham et al., 2023), these studies quantify susceptibility of the models’ predictions to changes in context or input. Such approaches are valid only if there is no direct causal link between the input and prediction that bypasses the explanation, which is rare in LMs (Bao et al., 2025). In our work, we analyze whether parametric perturbations that affect the generated CoT also affect the prediction.. The closest to ours are the contemporaneous works of Yeo et al. (2024) who use activation patching to measure causal effect of corrupting certain hidden states, and Zaman and Srivastava (2025) who use knowledge editing to evaluate existing (un)faithfulness metrics.
Background on Machine Unlearning.
Machine unlearning aims to remove some and only some undesired knowledge or behavior so as not to be regurgitated by LMs (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023). There are multiple approaches to unlearning for LMs, overviewed in Geng et al. (2025) and Appendix A. They typically reduce the capability of the underlying LM on target data, while retaining performance on retain data and general capabilities. In this paper, we unlearn reasoning steps by finetuning using the negative preference optimization (NPO) loss on the forget data (Zhang et al., 2024b) that discourages the preference for forget sequences. We add it to the KL divergence between the original and “unlearned” model’s predictions on the retain set Chen and Yang (2023); Yao et al. (2024). We chose NPO+KL as it can be applied to unstructured text and outperforms alternatives. More details in §4.1.
@@
Stage 1: InterventionStage 2: EvaluationUnlearning (§4.1)NPO+KL (Eq. 9)on individual CoT stepsControls (§4.2)1. Efficacy (Eq. 2) 2. Specificity (Eq. 3) 3. General capabilities $\Delta$prediction (§6.2)Prediction differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$.$\text{ff}_{\text{hard}}$ and $\text{ff}_{\text{soft}}$4.3)$\Delta$reasoning (§6.3)CoT Differencebetween $\mathcal{M}$ and $\mathcal{M}^{*}{}$
Figure 2: A high level overview of the two stages of pff: (1) parameter intervention and (2) evaluation. We instantiate pff with fur by using NPO+KL, controls to assure precision of unlearning and faithfulness metrics.
@@
A line of work has analyzed the sensitivity of models to perturbations applied to reasoning steps (Lanham et al., 2023; Bentham et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia) under the guise of faithfulness. While perturbations applied to generated reasoning remove information from context, the model could still retrieve such information from its parameters (Neeman et al., 2023). Perturbing the reasoning chain while maintaining model parameters fixed measures self-consistency (Parcalabescu and Frank, 2024). Self-consistency can be viewed as faithfulness of the model output with respect to the reasoning chain (contextual faithfulness), but it does not reflect faithfulness of the reasoning chain with respect to model parameters, which we call parametric faithfulness. Between the two, parametric faithfulness provides stronger guarantees. Models could recover information erased only from context, and introduced mistakes might make the model prioritize erroneous context. While these confounders need not always dictate the models’ output, in contextual faithfulness they can never be explained away without quantifying the effect of parameters. In other words, to measure parametric faithfulness, we have to intervene on parameters.
@@
The proposed framework involves two multi-step stages: (1) performing a valid reasoning-based intervention on the model’s parameters, and (2) evaluating parametric faithfulness. We outline our framework in Figure 2.
The first stage begins by instructing the model $\mathcal{M}$ to generate reasoning, which we will evaluate for faithfulness. The reasoning is broken into reasoning steps of a chosen granularity. Each individual reasoning step is used to guide an intervention on $\mathcal{M}$’s parameters, targeting those where a step’s information is stored. This produces a modified model, $\mathcal{M}^{*}$. Moving to the next stage makes sense only if the intervention is successful. Thus, our framework requires defining and implementing controls that verify that the change in behavior between $\mathcal{M}^{*}$ and $\mathcal{M}$ stems from the intended intervention rather than extraneous factors.
In the second stage, faithfulness is assessed with at least one of two evaluation protocols: (1) Instruct both $\mathcal{M}^{*}$ and $\mathcal{M}$ to directly give answers, then compute how often and how strongly their answers differ. (2) Instruct $\mathcal{M}^{*}$ and $\mathcal{M}$ to reason-then-answer, then compute how often they present different reasoning. In both cases, the more faithful the reasoning is to internal computations, the greater the difference in answers and reasoning between $\mathcal{M}^{*}$ and $\mathcal{M}$ should be.
@@ 4.1 Parameter Intervention
We erase knowledge contained in the verbalized reasoning steps using a preference-optimization unlearning method, NPO (Zhang et al., 2024b). Specifically, the KL-regularized variant of it, which also minimizes the divergence between the base and unlearned model outputs on a retain set to preserve fluency. [^2] We experimented with NPO+grad-diff, but results were always slightly worse than NPO+KL. We explored ROME and MEMIT (Meng et al., 2022, 2023), but they require a structured format, and do not perform well under paraphrases. We provide a detailed overview in Appendix A for readers unfamiliar with NPO+KL.
NPO+KL requires defining the forget set, $\mathcal{D}_{\text{FG}}$, and the retain set, $\mathcal{D}_{\text{RT}}$, which we construct as follows. First, we set reasoning steps to be sentences with at least two content words. For each step, we construct its $\mathcal{D}_{\text{FG}}$ of input-output pairs formed by taking, for each content word in the step, the prefix up to that word as input and the tokens of the content word as output. [^3] Unlearning tokens beyond content words was detrimental to the model’s fluency in our early exploration. NPO updates the model’s parameters to discourage it from predicting content words $\mathcal{D}_{\text{FG}}$ given prior context. We similarly construct $\mathcal{D}_{\text{RT}}$ for a given step as content words from four randomly selected CoT steps from other instances. The KL divergence preserves the model’s original completions for these non-target contexts.
We unlearn each reasoning step individually, for a total of $5$ iterations, and refer to the model obtained after unlearning the $i$-th reasoning step alone as $\mathcal{M}^{(i)^{*}}$. One unlearning iteration refers to a pass over $\mathcal{D}_{\text{FG}}$. We only update the second FF2 matrix of the Transformer MLPs, as this layer was found to act as a memory store (Geva et al., 2021b; Meng et al., 2022) and model editing methods frequently target it to update information (Meng et al., 2022, 2023; Hong et al., 2024). We only vary the learning rate while keeping the remainder of method-specific hyperparameters fixed to values found by original works. We report them in Appendix E.
@@ 4.2 Controls
$S=\frac{1}{\left|\mathcal{D}_{s}\right|}\sum_{k=1}^{\left|\mathcal{D}_{s}\right|}\mathds{1}[y_{k}=y^{*}_{k}].$ (3)
@@ 4.3 Faithfulness Measurements
where $r_{i}$ is the $i$-th reasoning step and $y^{(i)^{*}}$ the prediction made by $\mathcal{M}^{(i)^{*}}$ (after the $i$-th reasoning step is unlearned). [^5] A single faithful step is sufficient to show that the gist of the model’s internal reasoning is captured by the verbalized reasoning. In this sense, the entire chain can be considered faithful. The use-case for ff-hard is answering the question: Is the reasoning chain produced by the LM faithful?
@@
Base ARC-Challenge OpenbookQA Sports StrategyQA TruthfulQA
Model Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen Eff Spec Gen
LLaMA-8B $63.9$ $43.2$ $98.3$ $63.8$ $44.1$ $97.7$ $63.8$ $20.8$ $98.1$ $63.8$ $48.3$ $95.7$ $63.8$ $39.6$ $97.0$ $63.8$
LLaMA-3B $60.4$ $30.7$ $98.1$ $60.2$ $36.6$ $96.1$ $60.2$ $29.3$ $96.6$ $60.3$ $36.3$ $96.9$ $60.3$ $28.9$ $95.9$ $60.3$
Mistral-2 $59.0$ $71.5$ $96.4$ $58.9$ $72.1$ $97.6$ $58.8$ $50.6$ $94.8$ $59.0$ $65.4$ $96.3$ $59.0$ $48.6$ $95.0$ $59.0$
Phi-3 $69.9$ $40.8$ $99.5$ $69.6$ $44.2$ $99.4$ $69.6$ $31.1$ $97.0$ $69.9$ $18.7$ $98.2$ $69.9$ $11.0$ $97.4$ $69.8$
@@ 5 Experimental Setup
We employ five diverse multi-hop datasets: OpenbookQA (Book; Mihaylov et al., 2018), ARC-Challenge (Arc-ch; Clark et al., 2018), StrategyQA (SQA; Geva et al., 2021a), TruthfulQA (TQA; Lin et al., 2022) and the Sports understanding subtask of BigBench-Hard Srivastava et al. (2023). These datasets span a variety of domains, necessitating knowledge of science, sports, geography, health, law, finance and logic. We choose MCQA as the target task as it simplifies analysis of how the models’ predictive distribution shifts after unlearning due to availability of alternative answers. To retain comparable sizes, and due to expensive runtime of unlearning each CoT step, we select a subset of $250$ instances from the test split of each dataset to balance the question sources. [^6] For SQA, we use instances from the validation split due to the availability of labels. Sports has a total of $248$ instances. Details of datasets and models are in Appendix C.
@@
We use a two-step prompting approach Bowman et al. (2022); Lanham et al. (2023), where the model is first prompted to generate the CoT based on the question and answer options, and subsequently prompted to complete the answer letter based on the question, answer choices, and the CoT. We use greedy decoding when generating, producing a single CoT for each instance. For the prompts used, see Appendix D.
@@
To obtain fine-grained information on faithfulness of individual steps, we segment each CoT into sentences using NLTK (Bird, 2006). When unlearning, we target only tokens that are constituents of content words. [^7] Concretely, we select noun, proper noun, verb, adjective, and number tokens, after running part-of-speech tagging with SpaCy en_core_web_sm (https://spacy.io/). We opt for this approach so as to not unlearn the capability to verbalize reasoning from the models, but only knowledge within the steps, which we frequently observed prior to making this modification.
ARC-challenge OpenbookQA Sports StrategyQA TruthfulQA
Model fur +mistake fur +mistake fur +mistake fur +mistake fur +mistake
LLaMA-8B $\mathbf{39.6}$ $16.2$ $\mathbf{44.3}$ $18.0$ $29.3$ $30.0$ $30.7$ $\mathbf{32.3}$ $\mathbf{68.5}$ $25.0$
LLaMA-3B $\mathbf{64.4}$ $31.1$ $\mathbf{68.6}$ $45.9$ $64.9$ $65.5$ $\mathbf{71.0}$ $48.3$ $\mathbf{85.7}$ $32.9$
Mistral-2 $\mathbf{40.0}$ $31.6$ $\mathbf{60.0}$ $35.7$ $\mathbf{45.3}$ $36.8$ $\mathbf{48.2}$ $30.2$ 44.4 $30.3$
Phi-3 $\mathbf{39.1}$ $27.6$ $\mathbf{46.2}$ $38.5$ $\mathbf{54.0}$ $52.2$ $22.2$ $\mathbf{49.7}$ $29.1$ $\mathbf{31.9}$